# Gamma-ray burst Jet Progenitor: long vs short bursts by angular momentum and mass ratio

### v1

@@ 5.2 Homologous collapse of rigid rotating core
$\rho=\rho_{c}f^{3},\quad v_{s}^{2}=\gamma p/\rho=(v_{s}^{c})^{2}f^{3(\gamma-1)},$ (5.11)
@@
Here $\rho_{c}$ is the central density, and $v_{s}^{c}$ is the central sound velocity, given by $(v_{s}^{c})^{2}\equiv\gamma p_{c}/\rho_{c}$. The length scale function $a(t)$ relates to the central density and sound velocity by:
@@
Given the functional forms of $\rho$ and ${\bf u}$, the continuity equation (5.1) reduces to the trivial relation $\dot{f}=0$, indicating that the homologous profile $f(r)$ is time-independent. The integrated Euler equation (5.2) yields
@@
We are now in the position of postulating the rotation law for a rigidly and slowly rotating stellar core, which undergoes homologous gravitational collapse. Since the stellar rotating scale $\Omega^{-1}$ and collapsing timescales $\tau_{\rm grav}=R/\dot{R}$ are much larger than microscopic interaction and thermalization timescales, we approximately consider the gravitational collapse of rotating stellar core is an adiabatic process. Namely, at each collapse step at time $t$, the stellar core is approximately in an equilibrium state as if the collapsing stellar core adiabatically evolves in a time sequence of equilibrium states during gravitational collapse. The rigid core angular velocity $\Omega(t)$ increases in time. Therefore, we generalize the equilibrium rotation law (5.6) to
$j(t)=r^{2}\sin^{2}\theta a^{2}(t)\Omega(t);\quad\phi_{c}(t)=\frac{A^{2}(t)}{2}r^{2}\sin^{2}\theta a^{2}(t)\Omega(t),$ (5.15)
where $A(t)$ is the time-dependent rotation parameter.
To further describe the centrifugal potential $\phi_{c}$ of a slowly and rigidly rotating, homologously collapsing stellar core, we assume a time-dependent rotation parameter $A(t)$ in (5.15) depends on the centre sound velocity
$A^{2}(t)={\mathcal{A}}^{2}[v^{c}_{s}(t)]^{2},$ (5.16)
and ${\mathcal{A}}$ is a constant that does not depend on time. Thus, the centrifugal potential of a slowly and rigidly rotating and homologous collapsing core can be expressed by
$\phi_{c}=\frac{[v^{c}_{s}(t)]^{2}}{2}r^{2}\sin^{2}\theta{\mathcal{A}}^{2}a^{2}\Omega,\quad{\mathcal{A}}^{2}a^{2}\Omega={\mathcal{A}}^{2}a^{2}(t)\Omega(t),$ (5.17)
in the rotating frame. We will show soon that ${\mathcal{A}}a^{2}\Omega={\mathcal{A}}a^{2}(t)\Omega(t)$ is time-independent, based on total core mass and angular momentum conservation.
By integrating these equations, the evolution of the collapsing core can be numerically simulated or approximately analyzed, incorporating the effects of rotation and centrifugal potential. This approach can yield insights into the final state of the collapsing core, and help understand phenomena such as supernovae and gamma-ray bursts.
@@ 5.3 Spherical and homologous solution in space
Using the ansatz $v=(1/2)(\dot{a}/a)R^{2}$, Euler’s equation (5.14) is separable in spatial and temporal variables,
@@
where the positive eigenvalue $\lambda$ relates to the material binding energy of baryon fluid per energy density $\rho$ [16] or external pressure acting on the core surface. The core outer radius $r_{s}$ decreases as the $\lambda$ value increases. Equation (5.19) describes the time evolution of the length scale function $a(t)$ in the collapsing core.
The spatial solution to the Euler’s equation (5.18) is
@@
the boundary conditions are $f^{\prime}(0)=0$ and $f(0)=1$. Following [16] for the $\gamma=4/3$ solution, we obtain general solutions for $31<\gamma<4/$ 3 [17]. As shown in Fig. 1, three solutions of homologous density profile $f^{3}$ correspond to the selected values $(\gamma,\lambda)$:
@@
The homologous profile $f(r)$ becomes tangent to $f=0$ at the core outer radius $r_{s}$, namely $f^{\prime}(r_{s})\approx 0$. The core homologous profile $f(r)$ and outer radius $r_{s}$ do not depend on time.
In our analysis, we approximately decouple the centrifugal potential $\phi_{c}$ from the Euler equation (5.14), because the rotational potential energy is much smaller than the heat and gravitational potential energies. Namely, we neglect the back-reaction dynamics of the rotation potential $\phi_{c}$ in homologous gravitational collapse. The gravitational potential $\phi$ and heat function $h$ are balanced with the $\lambda$ term in Eq. 5.18. The homologously collapsing cores remain spherically symmetric without high moments.
@@
Figure 1: Colours on line for different eigenvalues (5.22). We show numerical solutions to the eigenvalue problem of Eqs. (5.18) and (5.19). The homologous density profile $f^{3}$ is plotted as a function of dimensionless radius $r=R/a$. The density profile $f^{3}$ does not change in time and vanishes at the core outer radius $r_{s}$ (5.22).
Figure 2: Corresponding to the density profile $f^{3}$ in Fig. 1, the scale length $a(t)/a_{\rm min}$ is plotted as in terms of $t/\tau_{\rm min}$, using units $a_{\rm min}$ (5.24) and $\tau_{\rm min}$ (5.25). Note that the initial time $t_{\rm initial}\sim 10^{7}\tau_{\rm min}\sim{\mathcal{O}}(1)$ second, given by the initial conditions (5.27). The shadow in between these colour lines indicates other possible physical solutions.
@@ 5.4 Initial central density and length scale function in time
The temporal equation (5.19) governs how the scale length $a(t)$ changes in time, and the solution is
$\frac{a(t)}{a_{\rm min}}$ $=$ $\left\{1+\Big{[}\frac{\lambda\gamma}{3(2-\gamma)(\gamma-1)^{2}}\Big{]}^{\frac{1}{2}}\Big{(}\frac{t}{\tau_{\rm min}}\Big{)}\right\}^{2-\gamma}.$ (5.23)
In a collapsing process from the initial time $t_{\rm intial}$ to the final time $t_{\rm final}=\tau_{\rm min}$, the scale function $a(t)$ decreases significantly, approaching the minimum,
@@
and $\tau_{\rm min}=(3\gamma)^{1/2}a_{\rm min}$. It gives the basic macroscopic length scale of the homologous collapse. The centre density $\rho_{c}$ and sound velocity $v_{s}^{c}$ increase in time,
@@
approaching their maximal values $\rho^{\rm max}_{c}=(3\gamma\kappa)^{-\frac{1}{\gamma-1}}$ and $1/\sqrt{3}$. The maximal centre density $\rho^{\rm max}_{c}=(3\gamma\kappa)^{-\frac{1}{\gamma-1}}$ is about $10$ times the nuclear saturation energy density $\rho_{0}\approx 2.4\times 10^{35}{\rm ergs}/{\rm cm}^{3}$, corresponding to the number density $n_{0}\approx 1.6\times 10^{38}/{\rm cm}^{3}$.
Instead of using collapsing time $t$, we choose the central density $\rho_{c}(t)$ as the primary variable to describe the homologous gravitational collapse of a stellar core. This choice is physically intuitive and aligns with the key stages of the collapse, characterized by changes in the core density. The initial value $\rho_{c}^{\rm in}=\rho_{c}(t_{\rm initial})$ depends on the homologous core mass $M$ and radius $r_{s}$. From Eqs. (5.23-5.26), we obtain
@@
for $t_{\rm initial}\sim 10^{6}~{}\tau_{\rm min}\sim{\mathcal{O}}(1)$ seconds. The final value $\rho_{c}^{\rm fi}=\rho_{c}(t_{\rm final})$ is the maximal value $\rho_{c}^{\rm max}$. The core undergoes a homologous collapsing process: starting from the initial time $t_{\rm initial}\sim{\mathcal{O}}(1)$ seconds and initial core centre density $\rho^{\rm in}_{c}$, ending at the final time $t_{\rm final}=\tau_{\rm min}\sim 10^{-5}$ seconds and final core centre density is $\rho^{\rm max}_{c}\approx 10\rho_{0}$. The duration is about ${\mathcal{O}}(1)$ seconds.
5.5 Core mass and angular momentum conservations
Using Eq. (5.21) and outer boundary condition $f^{\prime}(r_{s})\approx 0$ of the homologously collapsing core, we calculate its total core mass
@@ 5.5 Core mass and angular momentum conservations
$\approx$ $a^{5}(t)\rho_{c}(t)\Omega(t)\frac{\lambda\gamma}{4(\gamma-1)}\frac{8\pi}{15}r^{5}_{s}.$
@@
$a^{3}(t)\rho_{c}(t)={\rm const.},\quad a^{2}(t)\Omega(t)={\rm const.},$ (5.30)
@@
in the homologous collapsing process from the initial core density $\rho_{c}^{\rm in}$ to the final core density $\rho_{c}^{\rm fi}=\rho_{c}^{\rm max}$.
The ratio of total core mass and angular momentum
$J/M\approx(2/5)a^{2}(t)\Omega(t)r^{2}_{s},$ (5.31)
is constant in time, and proportional to the rotation parameter ${\mathcal{A}}^{2}a^{2}(t)\Omega(t)$ (5.17) for ${\mathcal{A}}^{2}=2r_{s}^{2}/5$. Therefore, we will use the ratio $J/M$ to characterize system rotation and mass energies. In our case, $J/M\ll 1$ for a slow-rotating system of the rotational energy being much smaller than mass-energy.
6 Photon-pair sphere/jet energetic properties
6.1 Photon-pair sphere/jet temperature profiles
@@
Based on the adiabatic approximation discussed in Sec. 5, we can use the “local” virial theorem (3.4) to obtain the baryon heat energy $dF$ (3.2) and temperature $T$ (3.4). As discussed in Sec. 4, the baryon heat energy is transformed into the thermal energy of a photon-pair plasma through hadronic photon production, resulting in a photon temperature $T_{\gamma}$ given by Eq. (4.3). Using these relations, the centrifugal potential (5.17) and homologous solution (5.20), we derive

### v2

@@ 5.3 Homologous collapse of rigid rotating core
$\rho=\rho_{c}(t)f^{3}(r),\quad v_{s}^{2}=\gamma p/\rho=(v_{s}^{c}(t))^{2}f^{3(\gamma-1)}(r),$ (5.12)
@@
Here $\rho_{c}(t)$ is the central density, and $v_{s}^{c}(t)$ is the central sound velocity, given by $(v_{s}^{c})^{2}\equiv\gamma p_{c}/\rho_{c}$. The length scale function $a(t)$ relates to the central density and sound velocity by:
@@
Given the functional forms of $\rho$ and ${\bf u}$, the continuity equation (5.1) reduces to the trivial relation $\dot{f}=0$, indicating that the homologous profile $f(r)$ is time-independent.
The integrated Euler equation (5.2) yields
@@
By integrating these equations, the evolution of the collapsing core can be numerically simulated or approximately analyzed, incorporating the effects of rotation and centrifugal potential.
@@ 5.4 Spherical and homologous solution in space
Neglecting $\phi_{c}$ and using the ansatz $v=(1/2)(\dot{a}/a)R^{2}$, Euler’s equation (5.15) is separable in spatial and temporal variables,
@@
where the positive eigenvalue $\lambda$ relates to the material binding energy of baryon fluid per energy density $\rho$ [29] or external pressure acting on the core surface. Namely, $\lambda dm$ relates to an elementary material binding energy for a baryon fluid element of mass $dm$ around the nuclear density. Equation (5.17) describes the time evolution of the length scale function $a(t)$ in the collapsing core. The time-independent homologous spatial solution to the Euler’s equation (5.16) is
@@
the boundary conditions are $f^{\prime}(0)=0$ and $f(0)=1$.
The mean thermal index $\gamma$ in the simple polytropic EoS (3.4), depending on the microscopic physics of material density and interactions, represents the average effect of repulsive heat energy pressure against attractive gravity force. Small $\gamma$ values for small heat pressure imply stellar cores would be unstable and undergo gravitationally collapses. Goldreich and Weber [29] show that such EoS (3.4) and $\gamma=4/3$ provide a reasonable approximation to collapsing stellar cores during the early phase before reaching nuclear density. In this simplified model, they obtain the homologous collapse solution of the onset value $\gamma=4/3$, below which massive cores undergo gravitational instability and collapse.
Following the same approach, we obtain [26] homologous collapse solutions for general values $31<\gamma<4/$ 3, studying how spherical fireballs form in gravitational collapses. However, the realistic situations of gravitational collapses of massive cores approaching nuclear density are much more complex due to nucleon interactions and possible phase transitions at different densities. These are beyond the scope of present studies. In this article, we approximately consider EoS (3.4) and the mean thermal index $\gamma\gtrsim 1$ as an effective $\mathcal{O}(1)$ free parameter relevant to how much the gravitational binding energy gains for forming fireballs. The larger (smaller) $\gamma$ value is, and the smaller (larger) gravitational binding energy gains. We select three $\gamma$ values (5.20) for illustration.
As shown in Fig. 1, three solutions of homologous density profile $f^{3}(r)$ correspond to the selected values $(\gamma,\lambda)$:
@@
The homologous profile $f(r)$ becomes tangent to $f=0$ at the core outer radius $r_{s}$, namely $f^{\prime}(r_{s})\approx 0$. The core homologous profile $f(r)$ and outer radius $r_{s}$ do not depend on time. The core outer radius $r_{s}$ decreases (increases) as the material binding energy $\lambda$ value increases (decreases).
In our analysis, we approximately decouple the centrifugal potential $\phi_{c}$ from the Euler equation (5.15), because the rotation energy is much smaller than the heat and gravitational potential energies. Namely, we neglect the back-reaction dynamics of the centrifugal potential $\phi_{c}$ in homologous gravitational collapse. The gravitational potential $\phi$ and heat function $h$ are balanced with the material binding energy $\lambda$ term in Eq. (5.16). The homologously collapsing cores remain spherically symmetric without high moments.
@@
Figure 1: Colours orange, blue and green on lines for three different eigenvalues selected in the list (5.20). We show numerical solutions to the eigenvalue problem of Eqs. (5.16) and (5.17). The homologous density profile $f^{3}(r)$ (5.12) is independent of the time. We plot it as a function of dimensionless radius $r=R/a$. It vanishes $f^{3}(r_{s})=0$ at the core outer radius $r_{s}$, see values in the list (5.20). The increase of the core density $\rho=\rho_{c}(t)f^{3}$ (5.12) in homologous collapse is due to the increase of the centre density $\rho_{c}(t)$ via the decrease of the length scale function $a(t)$ (5.14) and Fig. 2.
Figure 2: Corresponding to the density profile $f^{3}(r)$ in Fig. 1, the length scale $a(t)/a_{\rm min}$ is plotted as in terms of $t/\tau_{\rm min}$, using units $a_{\rm min}$ (5.22) and $\tau_{\rm min}$ (5.23). Note that the initial time $t_{\rm initial}\sim 10^{6}\tau_{\rm min}\sim{\mathcal{O}}(1)$ second, given by the initial conditions (5.25). Colours orange, blue and green on lines for three different eigenvalues indicated in the list (5.20). The shadow in between these colour lines indicates other possible physical solutions.
@@ 5.5 Initial central density and length scale function in time
In homologous collapses, the baryon core centre density $\rho_{c}(t)$ increases as the length scale $a(t)$ (5.14) decreases. The temporal equation (5.17) governs how the scale length $a(t)$ monotonically decreases in time. At the beginning of homologous collapses, i.e., $t_{\rm initial}$, the initial length scales and centre density $\rho^{\rm in}_{c}\equiv\rho_{c}(t_{\rm initial})$ are undetermined, varying from one homologous collapsing stellar system to others. In solving differential equation (5.17), we chose the second constant of time integration by using boundary conditions at the end $t_{\rm final}$ of homologous collapses, which give the minimal length scale $a_{\rm min}$ and the time scale $\tau_{\rm min}$. We use the $a_{\rm min}$ “sound horizon” and $\tau_{\rm min}$ as the basic units for the temporal variables $a(t)$ and $t$. In such a way, the solution is written in a compact form,
$\frac{a(t)}{a_{\rm min}}$ $=$ $\left\{1+\Big{[}\frac{\lambda\gamma}{3(2-\gamma)(\gamma-1)^{2}}\Big{]}^{\frac{1}{2}}\Big{(}\frac{t}{\tau_{\rm min}}\Big{)}\right\}^{2-\gamma},$ (5.21)
which describes a monotonically collapsing process inverse time from the initial time $t_{\rm intial}\not=0$ to the final time $t_{\rm final}=0$. The length scale function $a(t)$ decreases significantly, approaching the minimum,
@@
and $\tau_{\rm min}=(3\gamma)^{1/2}a_{\rm min}$. They give the basic macroscopic length and time scales of the homologous collapse. The centre density $\rho_{c}$ and sound velocity $v_{s}^{c}$ increase in time,
@@
approaching their maximal values $\rho^{\rm max}_{c}=(3\gamma\kappa)^{-\frac{1}{\gamma-1}}$ and $1/\sqrt{3}$. The maximal centre density $\rho^{\rm max}_{c}=(3\gamma\kappa)^{-\frac{1}{\gamma-1}}$ is about $10$ times the nuclear saturation energy density $\rho_{0}\approx 2.4\times 10^{35}{\rm ergs}/{\rm cm}^{3}$, corresponding to the number density $n_{0}\approx 1.6\times 10^{38}/{\rm cm}^{3}$. The reason for choosing the maximal baryon core central density $\rho^{\rm max}_{c}\approx 10\rho_{0}$ in this article is the following. We consider a massive stellar baryon core at nuclear saturation density $\rho_{0}$, about the density of neutron stars. The baryon core conserves its total mass $M$ and continuously collapses to a compact object of size about $2GM$ of black hole horizon. In this case, its average density increases and becomes about $5\sim 8$ times the nuclear saturation energy density.
Instead of using collapsing time $t$, we choose the core central density $\rho_{c}(t)$ as the primary variable to describe the homologous gravitational collapse of a stellar core. This choice is physically intuitive and aligns with the key stages of the collapse, characterized by changes in the core central density. The initial value $\rho_{c}^{\rm in}=\rho_{c}(t_{\rm initial})$ depends on the homologous core mass $M$ and radius $r_{s}$. From Eqs. (5.21-5.24), we obtain
@@
for $t_{\rm initial}\sim 10^{6}~{}\tau_{\rm min}\sim{\mathcal{O}}(1)$ seconds. The final value $\rho_{c}^{\rm fi}=\rho_{c}(t_{\rm final})$ is the maximal value $\rho_{c}^{\rm max}$. The core undergoes a homologous collapsing process: starting from the initial time $t_{\rm initial}\sim{\mathcal{O}}(1)$ seconds and initial core centre density $\rho^{\rm in}_{c}$, ending at the final time $t_{\rm final}=0$ and final core centre density is $\rho^{\rm max}_{c}\approx 10\rho_{0}$. The duration of homologous collapses $\Delta t=|t_{\rm final}-t_{\rm initial}|$ is about ${\mathcal{O}}(1)$ seconds.
5.6 The rotation law of homologous collapse core
We are now in the position of postulating the rotation law for a rigidly and slowly rotating stellar core, which undergoes homologous gravitational collapse. Since the stellar rotating scale $\Omega^{-1}$ and collapsing timescales $\tau_{\rm grav}=R/\dot{R}$ are much larger than microscopic interaction and thermalization timescales, we approximately consider the gravitational collapse of rotating stellar core is an adiabatic process. Namely, at each collapse step at time $t$, the stellar core is approximately in an equilibrium state as if the collapsing stellar core adiabatically evolves in a time sequence of equilibrium states during gravitational collapse.
The rigid core angular velocity $\Omega_{c}(t)$ monotonically increases in time, as gravitationally homologous collapses proceed. We generalize the equilibrium rotation law (5.9) to
$j(t)=r^{2}\sin^{2}\theta a^{2}(t)\Omega_{c}(t);\quad\phi_{c}(t)=\frac{[v^{c}_{s}(t)]^{2}}{2}r^{2}\sin^{2}\theta a^{2}(t)\Omega^{2}_{c}(t),$ (5.26)
and $\phi_{c}(t)=[v^{c}_{s}(t)]^{2}j(t)\Omega_{c}(t)/2$ in the rotating frame. We parameterize the time-varying centrifugal potential $\phi_{c}(t)$ proportional to $[v^{c}_{s}(t)]^{2}$, analogously to the gravitational potential $\phi$ and heat function $h$ (5.18) in this semi-analytical model.
We will show soon that the rotating parameter $a^{2}(t)\Omega_{c}(t)$ is time-independent, based on total core mass and angular momentum conservation.
5.7 Total core mass and angular momentum conservations
[…]
@@ 5.7 Total core mass and angular momentum conservations
$\approx$ $a^{5}(t)\rho_{c}(t)\Omega_{c}(t)\frac{\lambda\gamma}{4(\gamma-1)}\frac{8\pi}{15}r^{5}_{s},$
@@
$a^{3}(t)\rho_{c}(t)={\rm const.},\quad a^{2}(t)\Omega_{c}(t)={\rm const.},$ (5.29)
@@
in the homologous collapsing process from the initial core density $\rho_{c}^{\rm in}$ to the final core density $\rho_{c}^{\rm fi}=\rho_{c}^{\rm max}$. Equations (5.7) and (5.7) give the ratio of total core mass and angular momentum
$J/M\approx(2/5)a^{2}(t)\Omega_{c}(t)r^{2}_{s}.$ (5.30)
which is constant in time. This means the initial values from the beginning of collapses determine $J/M$ values. The larger $J/M$ values are, the larger rotating angular velocities $\Omega_{c}(t)$ are, and the $a^{2}\Omega^{2}_{c}$ values are large at the end of homologous collapses. The ratio $J/M$ is a physically sensible parameter not only for massive rotating core collapses, but also for massive rotating binary coalescence. Therefore, we will use the ratio $J/M$ to characterize rotating systems that undergo homologous collapses.
5.8 Slowly rotating approximation
Analogously, using the solution of gravitational potential $\phi$ (5.18) we calculate the total core gravitational binding energy
$U$ $=$ $\frac{1}{2}\int d^{3}R\rho\phi$ (5.31)
[…]
$\approx$ $-2M\frac{[v^{c}_{s}(t)]^{2}}{3(\gamma-1)}.$
Using the centrifugal potential (5.26) $\phi_{c}$, we calculate the total core rotation energy as
$U_{c}=\int d^{3}R\rho\phi_{c}=\frac{[v^{c}_{s}(t)]^{2}\Omega_{c}(t)}{2}\int d^{3}R\rho j=\frac{[v^{c}_{s}(t)]^{2}\Omega_{c}(t)}{2}J.$ (5.32)
It shows that as $\Omega_{c}(t)$ and $[v^{c}_{s}(t)]^{2}$ increase in homologous collapses, the core rotation energy $U_{c}$ increases by gaining gravitational binding energy. The ratio of total rotation energy and gravitational binding energy is
$U_{c}/|U|\approx\frac{3}{2}(\gamma-1)\Omega_{c}(t)(J/M).$ (5.33)
The rotation energy is much smaller than the gravitational binding energy when $\Omega_{c}(t)J/M\ll 1$ for slowly rotating cores in homologous collapses.
It must be emphasized that the theoretical framework is qualitatively valid only for $\Omega_{c}(t)J/M\ll 1$, since (i) we neglect the centrifugal potential $\phi_{c}$ in studying a homologous collapse in Secs. 5.4 and 5.5; (ii) we assume the rotation law (5.26) in homologous collapses. Nevertheless, we will see how the axial-symmetrically repulsive $\phi_{c}$ potential affects the spherical-symmetrically attractive potential $\phi$. The latter converts the gravitational binding energy into the photon-pair thermal energy of fireballs at the end of a homologous collapse [26]. We will then get an insight into the axial-symmetric distribution of photon-pair thermal energy (collimated fireball) when core rotations are present.
6 Collimated fireball energetic and geometric properties
@@
Based on the adiabatic approximation discussed in Sec. 5, we can use the “local” virial theorem (3.5) to obtain the baryon heat energy $dF$ (3.3) and temperature $T$ (3.5). As discussed in Sec. 4, the baryon heat energy is transformed into the thermal energy of a photon-pair plasma through hadronic photon production, resulting in a photon temperature (4.3).
6.1 Spherical fireball temperature of photons and $e^{+}e^{-}$ pairs
Figure 3: Colours orange, blue and green on line for three eigenvalues (5.20). Using results (6.1) and (6.2) for maximal sound velocity $v_{s}^{c}=1/\sqrt{3}$, we plot the homologous profiles of temperature $T_{\gamma}/m$ (left) and energy density $\rho_{\gamma}/\rho_{0}$ (right) as a function of dimensionless radius $r$. Note that the typical hadron mass $m\approx 1$ GeV and nuclear saturation density $\rho_{0}\approx 2.4\times 10^{35}{\rm ergs}/{\rm cm}^{3}$. These homologous profiles do not depend on the time $t$. But their absolute values depend on time-dependent sound velocity $(v_{s}^{c})^{2}(t)$, see Eqs. (6.1) and (6.2). The time evolutions of photon-pair sphere temperature $T_{\gamma}(r,t)/m$ (6.1) and energy density $\rho_{\gamma}(r,t)/\rho_{0}$ (6.2) are given in the Appendix of Ref. [26].
In non-rotating cases centrifugal potential $\phi_{c}=0$, we obtain the baryon temperature $T$ by using the local virial theorem (3.5) and homologous solution (5.18). The photon-pair temperature is given by $T_{\gamma}\approx 0.21T$ (4.3),
$\frac{T_{\gamma}}{m}$$\approx$ $0.21\frac{(v^{c}_{s})^{2}}{6}\left\{\frac{1}{\gamma(\gamma-1)}\left[(7\gamma-6)f^{3(\gamma-1)}(r)-\frac{\lambda}{6}\gamma r^{2}\right]\right\}.$ (6.1)
The time-dependent centre sound velocity $v^{c}_{s}(t)$ comes from Eq. (5.14). The temperatures $T=T(r,t)$ and $T_{\gamma}=T_{\gamma}(r,t)$ have the same spacetime configuration in homologous collapses. Moreover, we use photon productions (4.1) and (4.2) by baryon collisions to approximately obtain the photon-pair spherical fireball energy and number densities
$\frac{\rho_{\gamma}}{\rho_{0}}$$\approx$ $1.29\times\frac{4\alpha\alpha_{s}}{3\pi}\Big{(}\frac{T_{\gamma}}{m}\Big{)}^{4}\ln\left(1+\frac{2.9}{4\pi\alpha_{s}}\right),$ (6.2)
$\frac{n_{\gamma}}{n_{0}}$$\approx$ $6.16\times\frac{4\alpha\alpha_{s}}{3\pi}\Big{(}\frac{T_{\gamma}}{m}\Big{)}^{3}\ln\left(1+\frac{2.9}{4\pi\alpha_{s}}\right).$ (6.3)
In Fig. 3, we plot the photon-pair temperature $T_{\gamma}$ and energy density $\rho_{\gamma}$ at the end of homologous collapses when the centre sound velocity $v^{c}_{s}(t)$ approaches $1/\sqrt{3}$. The photon-pair temperature $T_{\gamma}$ reaches its maximum
$\frac{T_{\gamma}^{\rm max}}{m}\approx 0.21(v^{c}_{s})^{2}\frac{(7\gamma-6)}{6\gamma(\gamma-1)},$ (6.4)
at photon-pair sphere centre $r=0$. We define the outer boundary $r_{\gamma}$ of the photon-pair sphere (spherical fireball) by $T(r_{\gamma})=2m_{e}$, the energy threshold of electron-positron pair productions. Compared $r_{\gamma}$ with the baryon core boundary $r_{s}$ in Fig. 1, we find $r_{\gamma}<r_{s}$, namely the photon-pair sphere (fireball) is formed inside the baryon core, see also Fig. 5.
A posterior, as a self-consistency check, we find that the baryon temperature $T$ and photon temperature $T_{\gamma}$ are much smaller than baryon mass ($T_{\gamma}\approx 0.21T\ll m$). The condition is filled for applying the virial theorem (3.3).
Figure 4: Colours online. Using results (6.5) for maximal sound velocity $(v_{s}^{c})^{2}=1/3$, we plot the photon temperature $T_{\gamma}/m$ in the cylindrical coordinates $y=r\sin\theta$ and $z=r\cos\theta$, where the $\theta$ is the zenith angle. The colourmap scale (linear) is indicated by the colour bar $T_{\gamma}/m$ in the left, where the temperature unit is of typical baryon mass $m\approx 1$ GeV. The parameters are the ${\color[rgb]{0,0,1}blue}$ case parameters (5.20), and the initial conditions are (5.25). The left is a non-rotating case $a^{2}\Omega^{2}=0$. The right is a rotating case of $a^{2}\Omega^{2}=0.05$. The spherical fireball boundary $r_{\gamma}\lesssim 30$ and the collimated fireball boundary $r^{j}_{\gamma}(\theta)\leq r_{\gamma}$, see Fig. 5. Both are smaller than the baryon core outer radius $r_{s}\approx 36$, see the list (5.20) and blue curve in Fig. 1. The baryon temperature $T/m$ distribution in the cylindrical coordinates $y$ and $z$ is the same as the photon-pair temperature $T_{\gamma}=0.21T$.
6.2 Collimated fireball temperature of photons and $e^{+}e^{-}$ pairs
In rotating cases centrifugal potential $\phi_{c}\not=0$ (5.26), we obtain the baryon temperature $T$ by using the local virial theorem (3.5) and homologous solution (5.18). The photon-pair temperature is given by $T_{\gamma}\approx 0.21T$ (4.3),